# Evaluating the Impact Of Spatial Features Of Mobility Data and Index Choice On Database Performance

### v1

@@ 1. Introduction
Spatial data has become integral to modern applications due to the evergrowing amount of Internet-of-Things (IoT) devices. These devices produce geospatial information used in services such as real-time traffic-aware routing, incident hotspot detection, and weather updates and predictions, with various other use cases existing (Liebig et al., 2017; Karakaya et al., 2020, 2022; Mendelsohn et al., 2007). Building such applications requires informed decisions regarding the technology stack, especially regarding data storage factors.
@@
Choosing an optimal index and format for both spatial data and moving object data is non-trivial, as not all indexes are equally effective for all datasets and query types (Nam and Sussman, 2004). This is important to consider however, as the data format can impact the possible queries, the granularity of the query response, and the performance of the database. Performance may also vary depending on the characteristics of the data, such as spatial distribution and amount of data overlap, which can be difficult to determine a priori.
This paper investigates how spatial data characteristics, data format, and index choice impact database performance using PostGIS as a prototype database platform for our evaluation. We construct an application-driven benchmark using both synthetic and real-world spatial datasets with varying data distributions and degrees of overlap. We provide novel approximation methods to determine the degree of skew and overlap, which scale for large datasets by providing a constant time complexity using an approximation method. Our benchmark includes both read and write evaluations to fully regard the impact of dataset properties, index choice, and data format on database performance. Based on our results, we provide guidance for developers seeking to make informed storage decisions tailored to their use case.
@@
  • We develop novel, tunable approximation methods to assess overlap and distribution properties of trajectory-based datasets (Section 3).
@@ 2. Background and Related Work
In this section, we provide a background on moving object data, relevant storage and indexing strategies, and highlight key related work that guides our paper.
@@
Some databases provide interpolation features to support trajectory-based queries while storing data in a point-based format, with MobilityDB being a notable example (Zimányi et al., 2020).
@@
Spatial Indexes exist to support efficient querying of spatial data. These include R-Trees, QuadTrees, Generalized Search Trees (GiST), Block Range Indexes (BRIN), space-filling curves, and many more. Each of these have trade-offs in terms of data structure, dimensionality, and query performance. Many of these indexes rely on bounding boxes such as the Minimum Bounding Rectangle (MBR) to quickly filter irrelevant data points. A summary of common index types is shown in Table 1, however many more exist to also include other attributes such as time (Hadjieleftheriou et al., 2002; Tian et al., 2022). Systems exist that implement custom indexing and storage techniques for trajectory data  (Li et al., 2020; Cudre-Mauroux et al., 2010; Bakli et al., 2019).
Index Type R-Tree BRIN GiST SP-GiST Space-Filling Curves
@@
Benchmarking indexing performance requires realistic, diverse datasets. Using only one dataset, as is often done, fails to account for performance variations due to data characteristics
@@
Xu et al. proposed GMOBench, a benchmark for multi-modal trajectory data, but again relied on synthetic datasets and did not explore storage alternatives (Xu et al., 2015).
@@ 3. Approach
Real datasets, while sharing common characteristics, can differ greatly in terms of data skew and overlap. Data such as urban cycling data is often more evenly distributed across the city; however, when including all data from a country, a large skew may be present.
@@
Intersections refer to two vectors or trajectories intersecting at one or more points, while overlaps refer to these trajectories occupying the same space for a larger amount of distance besides a single point. A simple example would be a 4-way street crossing: Two bicycles going in perpendicular directions have intersecting paths, while two going in the same direction overlap. In spatial databases that use MBR-based indexing strategies however, these two terms come together due to the way indexes are structured. As mentioned in the previous section, MBRs help quickly filter irrelevant data when running queries and allow for faster query responses. When regarding two trajectories for a possible intersection, the system first checks these MBR structures for an overlap, leading to the interchangeable use of the two words in this context. Figure 2 shows how both actually intersecting and non-intersecting data can have overlapping MBRs.
@@
The amount of overlap in a dataset is crucial for the choice of the most suitable index, as some indexes perform better when data is space partitioned  (Aref and Ilyas, 2001). Given a trajectory t and another trajectory t’, we can determine if they overlap in matters of bounding box based-indexing by checking if their MBRs overlap. Using this, we can turn our trajectory data into a graph to use an existing graph metric to determine the amount of overlap in our dataset. We can represent each trajectory in our dataset as a node in a graph. If the MBR of two trajectories overlap, we create an edge between those nodes. An example of how the trajectory data is converted to a graph can be seen in Figure 3. A graph representation of our data enables us to apply graph density measures to quantify the trajectory overlap. The graph density sets the number of total edges in relation to the number of all possible edges. Given a set of nodes $V$ and edges $E$, the density of a graph can be calculated as follows:
@@
Going back to our way of converting trajectories to overlaps, density in our case reflects the amount of overlapping trajectories in relation to the total amount of possible overlaps. A high density (i.e closer to 1) indicates that a high degree of trajectories overlap with one another, while a low value (closer to 0) indicates the opposite. The general graph approach is not without its problems, especially with highly overlapping data: Generating a graph data structure from large trajctory datasets, containing millions or billions of instances, is computationally not feasible. We thus propose an approximation to allow developers a fast estimation of their dataset characteristics. Given a dataset of m trajectories:
@@
This allows for a constant time complexity of $O(n*p)$ instead of $O(m^{2})$ for all graph sizes, where we can choose $n$ and $p$ based on the desired accuracy of our approximation. In the rest of the paper, we refer to this metric as the global overlap coefficient (GOC) of a dataset.
@@
A value of $<1$ indicates a clustered distribution, while a value of $>1$ indicates a distributed dataset. The larger the value, the more distributed the dataset is. When applying this to trajectories, the idea does not hold up well: The distance between trajectories is the closest possible distance of the two, which is not indicative of their actual distribution. Two trajectories may intersect at one point, but be very far apart otherwise and still have a distance of 0. We can however still use the ANN idea if we convert our trajectories into multiple points. In this paper, we rely on the update frequency of the trajectory to convert trajectories, meaning that every time a sensor updates its position, we create a point in our dataset.
Figure 4. These simplified trajectories show how we can apply ANN to trajectories. With a small number of trajectories, exactly calculating this value is still realistic. Including a large amount of trajectories necesitates an approximation to finish the calculation in a reasonable time. Our ANN approximation excludes points from the own trajectory.
@@
  • We scale this value by multiplying it with $(m*k)/n$ to get an approximation of $D_{O}$.
@@
This allows us to reduce the time complexity to $O(n*p)$, where we can choose $n$ and $p$ based on the desired accuracy of our approximation instead $O((m*k)^{2})$ for all dataset sizes.
@@
4. Evaluation
Figure 5. We include 7 different datasets in our evaluation, with 4 synthetic and 3 real-world datasets. The real-world datasets are from the SimRa project, the Deutsche Flugsicherung, and the Piraeus AIS dataset.
@@ 4. Evaluation
Random (Random) Random trajectories 3,000,000 0.0032 27.053
Evenly Distributed (Even) Evenly distributed trajectories 3,000,000 0.0036 27.310
@@
SimRa (Cycling) Cycling trajectory data from Berlin 4,387 0.106 5.052
Deutsche Flugsicherung (Aviation) Flight trajectory data from Northrhine-Westphalia 215,908 0.201 0.698
Piraeus AIS Dataset (AIS) AIS ship data from a research maritime dataset. 2,719 0.392 0.996
@@ 4.1. Experiment Design
Each of these queries is run with 50 unique configurations. Using a Contains and a single Insert query as an example: During each of the configurations, the bounding box of the polygon in which trajectories are queried is unique and a randomly generated trajectory is inserted into the database within a specified polygon. The SUT is therefore subjected to 50 different benchmark configurations for each query type. We take further considerations into account here, such as filtering polygons which return no results using rejection sampling, and ensuring that the bounding box is within the area of the dataset. Each of these experiments is run against each dataset in every unique combination of index type and data format that we include in this paper. Each dataset is included in two formats, a segmented and non-segmented version, and three different index types are included in the evaluation (GiST, SP-GiST, and BRIN). As SP-GiST is specifically designed for space partitioned data, our assumption is that datasets with low overlap will benefit from this index type. The segmented version of the dataset is created by splitting the trajectory using the update frequency of the dataset.
@@ 4.3.1. Read Performance
(a) Segmented data
(b) Non-segmented data
Figure 6. On average, GiST outperforms SP-GiST, especially in the segmented formats, however there are scenarios where SP-GiST does provide an advantage. If designing a general purpose application able to handle a variety of queries, GiST is the better choice. BRIN performed poorly across all datasets and query types, and should likely not be used for trajectory data.
@@
This lead us to further investigate the relation between GOC and the average speedup from relying on the non-segmented format when using GiST as an index type when averaging across all query types. Figure 8 shows the relation between GOC and average speedup across all query types when using GiST as an index type, with the dotted line indicating a possible regression between values.
Figure 8. In our evaluation, datasets with a lower GOC benefitted heavily from using the non-segmented format when averaging across all query types. Higher GOC datasets received little to no speedup, with AIS data even performing better when using the segmented format. We were not able to determine a significant correlation however.
@@ 4.3.2. Write Performance
Figure 9 shows the performance of the different index types across all datasets and query types, with the singular operations on the left and the batch operations on the right. Our results show that BRIN is the best performing index type for most write operations, however the benefit is not as pronounced as it was negative for read operations. As its benefit are inconsistently spread across write operations, we believe that it is not a good choice for trajectory data in PostGIS unless an application is heavily write focused. SP-GiST was outperformed by GiST in nearly all of our cases, however it sometimes provides a small advantage in datasets with low overlap. In single insert scenarios, SP-GiST was able to outperform GiST on average in all datasets, however the performance difference was negligible.
@@
When regarding the impact of data format on write performance, we found that the segmented format lacked behind the non-segmented trajectory in nearly all of our comparisons. This is as expected, due to the fact that each operation on a segmented trajectory requires the database to perform the operation on multiple rows, whether it be inserting a trajectory in a segmented format or deleting/updating an existing one. The difference is not as pronounced in our insert operations, as we do not insert a percentage of the dataset, but a fixed number of trajectories.
@@ 4.4. Summary of Findings & Recommendations for Developers
While we still believe that trajectory ANN to be an important metric, it did not impact performance in our evaluation. When relying on another indexing strategy and database, it may be beneficial to include ANN in the evaluation again to regard a possible correlation with performance.
@@
When implementing a segmented data format, both index types perform similarly, with SP-GiST performing slightly better in some cases. These cases are however not tied to the degree of overlap in the data.
Our findings show that the higher the GOC of your data, the less of a benefit it is to store data in a non-segmented format. When using GiST, High GOC datasets benefit only slightly and in some cases even suffer from the non-segmented format. For those datasets, we recommend using the segmented format, as it provides better performance and a higher level of detail.
While adapting ANN to trajectories provides novel insights into a dataset, we did not find a correlation between ANN and performance. While other experiments may show a relationship here, our results indicate that data skew does not impact performance when using trajectory-based data.
When running a write-heavy workload, the non-segmented format provides advantage as it allows for faster writes across all datasets. While BRIN indexes are a good choice if one only considers write performance, its poor read performance likely makes it unsuitable for most applications.
If expecting a mix of read and write queries, we recommend evaluating the GOC of your data to determine the optimal data format for performance, and relying on GiST as the index type. Our assumption was that SP-GIST would prove a better choice in scenarios where data is equally distributed across the observed area, as it was specifically designed for cases where data is spatially partitioned. However, results show that at least for trajectory data in PostGIS, this is not the case. Regarding write performance, both index types perform similarly in most cases, with BRIN having an advantage in most scenarios. The GOC seems to be a good indicator whether the non-segmented format has any benefit at all, while the ANN coefficient did not seem to impact performance in our evaluation.

### v2

@@ 1. Introduction
Spatial data has become integral to modern applications due to the ever-growing amount of Internet-of-Things (IoT) devices. These devices produce geospatial information used in services such as real-time traffic-aware routing, incident hotspot detection, and weather updates and predictions, with various other use cases existing (Liebig et al., 2017; Karakaya et al., 2020, 2022; Mendelsohn et al., 2007). Building such applications requires informed decisions regarding the technology stack, especially regarding data storage factors.
@@
Choosing an optimal index and format for both spatial data and moving object data is non-trivial, as not all indexes are equally effective for all datasets and query types (Nam and Sussman, 2004). This is important to consider, as the data format can impact the possible queries, the granularity of the query response, and the performance of the database. Performance may also vary depending on the characteristics of the data, such as spatial distribution and amount of data overlap, which can be difficult to determine in advance.
This paper investigates how spatial data characteristics, data format, and index choice impact database performance using PostGIS as a prototype database platform for our evaluation. We construct an application-driven benchmark using both synthetic and real-world spatial datasets with varying data distributions and degrees of overlap. We provide novel approximation methods to determine the degree of skew and overlap, which scale for large datasets by introducing a Monte Carlo method which calculates the degrees within 3% of their respective values using less than 1% of the required time it takes to accurately calculate them. Our benchmark includes both read and write evaluations to fully regard the impact of dataset properties, index choice, and data format on database performance. Based on our results, we provide guidance for developers seeking to make informed storage decisions tailored to their use case.
@@
  • We develop novel metrics and tunable approximation methods to assess overlap and distribution properties of trajectory-based datasets (Section 3).
@@ 2. Background and Related Work
Index Type R-Tree BRIN GiST SP-GiST Space-Filling Curves
In this section, we provide a background on moving object data, relevant storage and indexing strategies. Lastly, we highlight key related work that guides our paper.
@@
Some databases provide interpolation features to support trajectory-based queries while storing data in a point-based format, with MobilityDB being a notable example (Zimányi et al., 2020). Here, different interpolation strategies can be applied, with linear interpolation being the most common.
@@
Spatial Indexes exist to support efficient querying of spatial data. These include R-Trees, QuadTrees, Generalized Search Trees (GiST), Block Range Indexes (BRIN), space-filling curves, and many more. Each of these have trade-offs in terms of data structure, dimensionality, and query performance. Many of these indexes rely on bounding boxes such as the Minimum Bounding Rectangle (MBR) to quickly filter irrelevant data points. A summary of common index types is shown in Table 1, however, many more exist to also include other attributes such as time (Hadjieleftheriou et al., 2002; Tian et al., 2022). Systems exist that implement custom indexing and storage techniques for trajectory data  (Li et al., 2020; Cudre-Mauroux et al., 2010; Bakli et al., 2019).
@@
Benchmarking indexing performance requires realistic, diverse datasets. Using only one dataset, as is often done, fails to account for performance variations due to data characteristics.
@@
Xu et al. proposed GMOBench, a benchmark for multi-modal trajectory data, but again relied on synthetic datasets and did not explore storage alternatives (Xu et al., 2015). This approach however did consider the possible impact of storing 3 or 4-dimensional trajectory data.
@@ 3. Approach
Real datasets, while sharing common characteristics, can differ greatly in terms of data skew and overlap. Data such as urban cycling data is often more evenly distributed across the city, however, when including all data from a country, a large skew may be present.
@@
Intersections refer to two vectors or trajectories intersecting at one or more points, while overlaps refer to these trajectories occupying the same space for a larger amount of distance besides a single point. A simple example would be a 4-way street crossing: Two bicycles going in perpendicular directions have intersecting paths, while two going in the same direction overlap. In spatial databases that use MBR-based indexing strategies however, these two terms come together due to the way indexes are structured. As mentioned in the previous section, MBRs help quickly filter irrelevant data when running queries and allow for faster query responses. When regarding two trajectories for a possible intersection, the system first checks these MBR structures for an overlap, leading to the interchangeable use of the two words in this context. Figure 2 shows how non-intersecting data can have overlapping MBRs.
@@
The amount of overlap in a dataset is crucial for the choice of the most suitable index, as some indexes perform better when data is space partitioned  (Aref and Ilyas, 2001). Given a trajectory t and another trajectory t’, we can determine if they overlap in matters of bounding-box-based-indexing by checking if their MBRs overlap. Using this, we can turn our trajectory data into a graph to use an existing graph metric to determine the amount of overlap in our dataset. We can represent each trajectory in our dataset as a node in a graph. If the MBR of two trajectories overlap, we create an edge between those nodes. An example of how the trajectory data is converted to a graph can be seen in Figure 3. A graph representation of our data enables us to apply graph density measures to quantify the trajectory overlap. The graph density sets the number of total edges in relation to the number of all possible edges. Given a set of nodes $V$ and edges $E$, the density of a graph can be calculated as follows:
@@
Going back to our way of converting trajectories to overlaps, density in our case reflects the amount of overlapping trajectories in relation to the total amount of possible overlaps, with its value ranging from 0 to 1. A high density (i.e closer to 1) indicates that a high degree of trajectories overlap with one another, while a low value (closer to 0) indicates the opposite. The general graph approach is not without its problems, especially with highly overlapping data: Generating a graph data structure from large trajctory datasets, containing millions or billions of instances, is computationally not feasible. We thus propose an approximation to allow developers a faster estimation of their dataset characteristics. Given a dataset of m trajectories:
@@
This allows for a time complexity of $O(n^{2}\cdot p)$ instead of $O(m^{2})$ for all graph sizes, where we can choose $n$ and $p$ based on the desired accuracy of our approximation. Logically, the higher we set $n$ and $p$, the more accurate our approximation will be, at the cost of longer computation time. In the rest of the paper, we refer to this metric as the global overlap coefficient (GOC) of a dataset.
Calculating the exact GOC of a dataset, while possible, is often not feasible in a reasonable time frame. Table 2 shows the results and calculation time of our GOC approximation method, highlighting the noticeable increase in calculation time with increasing $n$ and $p$ values. The highest approximation that we include ($n$ = 100, $p$ = 10,000) took over 2 hours to calculate, while the exact value took noticeably longer. We recommend to not accurately calculate dataset GOC due to the time requirement, considering the likely sufficient approximation results.
@@
A value of <1<1< 1 indicates a clustered distribution, while a value of >1>1> 1 indicates a distributed dataset. The larger the value, the more distributed the dataset is. When applying this to trajectories, the idea does not hold up well: The distance between trajectories is the closest possible distance of the two, which is not indicative of their actual distribution. Two trajectories may intersect at one point, but be very far apart otherwise and still have a distance of 0. We can however still use the ANN idea if we convert our trajectories into multiple points. In this paper, we rely on the update frequency of the trajectory to convert trajectories, meaning that every time a sensor updates its position, we create a point in our dataset.
Figure 4. These simplified trajectories show how we can apply ANN to trajectories. With a small number of trajectories, exactly calculating this value is still realistic. Including a large amount of trajectories necessitates an approximation to finish the calculation in a reasonable time. Our ANN approximation excludes points from the own trajectory.
Table 2. We use the aviation dataset also included in the evaluation to evaluate the accuracy of our approximation methods. Our approximations show that even with a small $n$ and $p$, we can achieve a near perfect approximation of the GOC and ANN values in a fraction of the required calculation time.
(a) GOC
$n$ $p$ GOC Calculation Time (s)
10 100 0.3759 4.25
10 500 0.4083 20.64
10 1,000 0.4017 50.22
10 10,000 0.4025 625.85
100 100 0.3985 51.70
100 500 0.3995 233.20
100 1,000 0.4005 548.87
100 10,000 0.4012 7525.20 ($\sim$2h)
1 215,908 0.4012 96293.23 ($\sim$27h)
(b) ANN
$n$ $p$ ANN Calculation Time (s)
10 100 0.7002 2.65
10 500 0.6982 4.55
10 1,000 0.6943 11.19
10 10,000 0.6952 100.95
100 100 0.6908 9.1
100 500 0.6960 45.78
100 1,000 0.6954 92.3
100 10,000 0.6942 918.96
1 $\sim$30M 0.6942 28621.15 ($\sim$8h)
@@
  • We scale this value by multiplying it with $(m\cdot k)/n$ to get an approximation of $D_{O}$ and calculate $D_{E}$ using the same formula as before (using the original amount of points in the dataset).
@@
This allows us to reduce the time complexity to $O(n^{2}\cdot p)$, where we can choose $n$ and $p$ based on the desired accuracy of our approximation instead $O((m\cdot k)^{2})$ for all dataset sizes. We exclude all points from the own trajectory, as including these would not provide any valuable information regarding the dataset when evaluating trajectories. We evaluate the accuracy of our approximations of GOC and ANN by comparing the approximated values to the exact results, using different $n$ and $p$ values to provide an overview of the accuracy of our methods. Table 2 shows the results of our approximation method. Calculating the exact ANN value for a dataset containing $\sim$ 30 million points took over 8 hours, while our approximation method was able to achieve a near perfect approximation in a fraction of the time. Assuming even larger datasets, we consider an approximation to be necessary to provide a reasonable time frame for developers to evaluate their dataset characteristics.
@@
Figure 5. We include 7 different datasets in our evaluation, with 4 synthetic and 3 real-world datasets. The real-world datasets are from the SimRa project [^1] https://github.com/simra-project/dataset, the Deutsche Flugsicherung [^2] The dataset is not open-source, but can be requested from the Deutsche Flugsicherung., and the Piraeus AIS dataset [^3] https://zenodo.org/records/6323416.
@@
Random (Random) Random trajectories 3,000,000 0.0064 27.053
Evenly Distributed (Even) Evenly distributed trajectories 3,000,000 0.0072 27.310
@@
SimRa (Cycling) Cycling trajectory data from Berlin and Potsdam 4,387 0.212 5.052
Deutsche Flugsicherung (Aviation) Flight trajectory data from Northrhine-Westphalia 215,908 0.4012 0.694
Piraeus AIS Dataset (AIS) AIS ship data from a research maritime dataset 2,719 0.784 0.996
4. Evaluation
@@ 4.1. Experiment Design
(a) Segmented data
(b) Non-segmented data
Figure 6. On average, GiST outperforms SP-GiST, especially in the segmented formats, however, there are scenarios where SP-GiST does provide an advantage. If designing a general purpose application able to handle a variety of queries, GiST is the better choice. BRIN performed poorly across all datasets and query types, and should likely not be used for trajectory data.
@@
Each of these queries is run with 50 unique configurations. Using a Contains and a single Insert query as an example: During each of the configurations, the bounding box of the polygon in which trajectories are queried is unique and a randomly generated trajectory is inserted into the database within a specified polygon. The system under test (SUT) is therefore subjected to 50 different benchmark configurations for each query type. We take further considerations into account here, such as filtering polygons which return no results using rejection sampling, and ensuring that the bounding box is within the area of the dataset. Each of these experiments is run against each dataset in every unique combination of index type and data format that we include in this paper. Each dataset is included in two formats, a segmented and non-segmented version, and three different index types are included in the evaluation (GiST, SP-GiST, and BRIN). As SP-GiST is specifically designed for space partitioned data, our expectation is that datasets with low overlap will benefit from this index type. The segmented version of the dataset is created by splitting the trajectory using the update frequency of the dataset.
@@ 4.3.1. Read Performance
This lead us to further investigate the relation between GOC and the average speedup from relying on the non-segmented format when using GiST as an index type when averaging across all query types. Figure 8 shows the relation between GOC and average speedup across all query types when using GiST as an index type, with the dotted line indicating a possible interpolation between values.
Figure 8. In our evaluation, datasets with a lower GOC benefitted heavily from using the non-segmented format when averaging across all query types. Higher GOC datasets received little to no speedup, with AIS data even performing better when using the segmented format. We were not able to determine a significant correlation however ($p=0.0564,\alpha=0.05$).
@@ 4.3.2. Write Performance
Figure 9 shows the performance of the different index types across all datasets and query types, with the singular operations on the left and the batch operations on the right. Our results show that BRIN is the best performing index type for most write operations, however, the benefit is not as pronounced as it was negative for read operations. As its benefits are inconsistently spread across write operations, we believe that it is not a good choice for trajectory data in PostGIS unless an application is heavily write focused. SP-GiST was outperformed by GiST in nearly all of our cases, however, it sometimes provides a small advantage in datasets with low overlap. In single insert scenarios, SP-GiST was able to outperform GiST on average in all datasets, however, the performance difference was negligible.
@@
When regarding the impact of data format on write performance, we found that the segmented format lagged behind the non-segmented trajectory in nearly all of our comparisons. This is as expected, due to the fact that each operation on a segmented trajectory requires the database to perform the operation on multiple rows, whether it be inserting a trajectory in a segmented format or deleting/updating an existing one. The difference is not as pronounced in our insert operations, as we do not insert a percentage of the dataset, but a fixed number of trajectories.
@@ 4.4. Summary of Findings & Recommendations for Developers
Figure 10. Unless the expected workload is extremely write-heavy, relying on BRIN is not recommended. The shaded area around the mean indicates the 95% confidence interval, and the yellow line is nearly identical to the light blue one. The mean and interval are calculated across all datasets and query types.
@@
While we still believe trajectory ANN to be an important metric, it did not impact performance in our evaluation. When relying on another indexing strategy and database, it may be beneficial to include ANN in the evaluation again to regard a possible correlation with performance.
@@
When implementing a segmented data format, both index types perform similarly, with SP-GiST performing slightly better in some cases. These cases are however, not tied to the degree of overlap in the data.
Our findings show that the higher the GOC of the data, the less of a benefit it is to store data in a non-segmented format. When using GiST, high GOC datasets benefit only slightly and in some cases even suffer from the non-segmented format. For those datasets, we recommend using the segmented format, as it provides better performance and a higher level of detail. While adapting ANN to trajectories provides novel insights into a dataset, we did not find a correlation between ANN and performance. While other experiments may show a relationship here, our results indicate that data skew does not impact performance when using trajectory-based data in PostGIS. ANN however, could be interesting to regard when evaluating other databases and/or indexing strategies present in such database systems.
When running a write-heavy workload, the non-segmented format provides an advantage as it allows for faster writes across all datasets. While BRIN indexes are a good choice if one only considers write performance, its poor read performance likely makes it unsuitable for most applications. BRIN may show better read performance in scenarios where data is inserted in a sorted order, i.e, with an always increasing timestamp, and then running time-based queries. However, as such data is not trajectory-based, we did not include it in our evaluation.
If expecting a mix of read and write queries, we recommend evaluating the GOC of the data to determine the optimal data format for performance, and relying on GiST as the index type. We recommend to also keep the level of required detail in mind, and to regard if the non-segmented format is at all beneficial or usable for the intended application. Our expectation was that SP-GiST would prove a better choice in scenarios where data is equally distributed across the observed area, as it was specifically designed for cases where data is spatially partitioned. However, results show that at least for trajectory data in PostGIS, this is not the case. Regarding write performance, both index types perform similarly in most cases, with BRIN having an advantage in most scenarios. As previously mentioned, the downside of BRIN is in its read performance, which likely does not balance out the performance benefits even in write-heavy workloads.
In Figure 10, we show that if one’s application is not extremely write-heavy, the performance benefits of BRIN do not outweigh the read performance benefits of GiST and SP-GiST. Even in the best case scenario (5%/95% Read/Write Ratio) for BRIN that we evaluated, the mean operation time is still lower for GiST in both cases and for SP-GiST in the non-segmented format.
The GOC seems to be a good indicator whether the non-segmented format has any benefit at all, while the ANN coefficient did not seem to impact performance in our evaluation. We assume that other indexing methods and datasets outside of the evaluated ones may show a correlation between ANN and performance, specifically those which are designed for non-skewed data.